# Examining dark neutron decay using f-mode gravitational waves

### v2

@@ Abstract
These findings highlight the significance of gravitational wave detections in investigating neutron star structure, particularly with next-generation observatories such as the Einstein Telescope and Cosmic Explorer. In particular, the f-mode frequencies generated by such stars fall within the detection range of modern telescopes, making it possible to observe these signals directly. This opens a promising pathway for probing dark matter indirectly—using neutron stars as natural laboratories for fundamental physics.
@@ 1 Introduction
The fundamental mode (f-mode) is a type of non-radial oscillation in compact stars, primarily governed by the star’s bulk properties—such as mass, radius, and internal composition. These oscillations generate gravitational waves, and their frequency and damping time are sensitive to the equation of state (EoS) of the stellar interior. As such, the f-mode serves as a powerful diagnostic tool for probing the internal structure and composition of neutron stars and, more generally, for constraining the nature of dense matter under extreme conditions.
@@
Furthermore, the damping time and frequency of the f-mode are also linked to the star’s properties, such as mass, compactness, and moment of inertia, quantities that could be inferred from spin-orbit coupling effects in binary pulsar systems [27]. These interdependencies allow the f-mode to act as a bridge between observable gravitational wave signatures and the otherwise inaccessible interior physics of compact stars.
In this study, the neutron decay into dark matter particle model is tested using well-established universal relations for neutron stars (NSs) involving f-mode properties—relations that have been foundational in the field due to their near-independence from the specific choice of nuclear equation of state (EoS), and which have recently been revisited [28, 29]. This investigation is further extended to isolated compact stars composed of dark matter (DM), modeled as static, spherically symmetric objects with EoSs for both fermionic and bosonic dark matter particles.
This paper has been divided into different sections. The layout of this study is as follows: Section 2 provides theoretical possibility of neutron decay into dark matter particles, examining potential mechanisms and astrophysical constraints. Section 3 provides a comprehensive overview of neutron stars, detailing their formation, key properties, and observed phenomena. Section 4 focuses on the crucial concept of the Equation of State (EoS) governing the ultradense matter within neutron stars, exploring theoretical models and their implications. Building upon this, Section 5 discusses the fundamental structural equations that describe the equilibrium of neutron stars under their immense gravity. Section 6 further elaborates on the Equation of State, potentially exploring specific aspects or advanced considerations. Section 7 analyzes the intriguing phenomenon of universalities observed in neutron star properties, highlighting equation-of-state independent relationships. This section will be further subdivided, with Section 7.1 specifically focusing on the damping time of neutron star oscillations and its physical mechanisms. Section 7.2 will then examine the mass-scaled angular frequency of rotating neutron stars and its significance. Section 8 provides a comparative analysis of f-mode frequencies across various neutron star models, offering insights into their structural and dynamical characteristics. Section 9 concludes the article by summarizing the key findings and discussing their broader relevance to neutron star physics and gravitational wave astronomy.
@@ 4 Equation of State
$\epsilon_{H}=\frac{1}{2}m^{2}_{\sigma}\sigma^{2}+\frac{1}{2}m^{2}_{\omega}\omega^{2}+\frac{1}{2}m^{2}_{\rho}\rho^{2}+\frac{1}{\pi^{2}}\int^{k_{F}^{n}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{p}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk$ (4.2)
$+\frac{1}{\pi^{2}}\int^{k_{F}^{e}}_{0}k^{2}\sqrt{k^{2}+m_{e}^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{\mu}}_{0}k^{2}\sqrt{k^{2}+m_{\mu}^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{\chi}}_{0}k^{2}\sqrt{k^{2}+m_{\chi}^{2}}dk\,.$
where $m_{i}$ is the mass of corresponding particle (subscripted). The nucleon effective mass, $M_{N}^{*}(\sigma)$, given by [108, 109]
$M_{N}^{*}(\sigma)=M_{N}-g_{\sigma}\sigma+\frac{d\tilde{(}g_{\sigma}\sigma)^{2}}{2}\,.$ (4.3)
Here $g_{\sigma}$ is the $\sigma$-nucleon coupling constant in free space calculated in the model. The pressure ($P$) is calculated using
$P=\int\mu_{f}n_{f}-\epsilon.$ (4.4)
The nucleons-only EoS is taken from [108] and strange matter EoS is drive based on MIT bag model and the EoS used is given by
$P=\frac{1}{3}(\epsilon-4B)$ (4.5)
where $\epsilon$ the energy density of the strange matter and $B$ is the bag constant, taken to be 1014gm/fm3 [113].
@@ 5 Structural equations
Neutron stars are dense astrophysical objects shaped by intense nuclear and gravitational forces. The potential presence of dark matter within these stars may influence their structural characteristics, resulting in changes to their equation of state (EoS). This study examines methods for integrating dark neutron components into the conventional neutron star EoS and solving the Tolman-Oppenheimer-Volkoff (TOV) [114, 115] equations to assess their impact
@@
$\frac{dP}{dr}=-\frac{(\epsilon+P)\frac{d\Phi}{dr}}{r}$ (5.1)
$\frac{d\Phi}{dr}=\frac{M+4\pi r^{3}P}{r(r-2M)}$ (5.2)
$\frac{dM}{dr}=4\pi r^{2}\epsilon$ (5.3)
where $P$ is the pressure, $\epsilon$ is the energy ensity, $\Phi$ is the gravitational potential, $M$ is the mass, and $r$ is the radial coordinate
@@
The f-mode oscillations are governed by two coupled differential equations [116, 117, 118]:
$\frac{dW}{dr}=\frac{d\epsilon}{dP}\left[\omega^{2}r^{2}e^{\lambda}-2\Phi V+\frac{d\Phi}{dr}W\right]-l(l+1)e^{\lambda}V$ (5.4)
$\frac{dV}{dr}=2\frac{d\Phi}{dr}V-e^{\lambda}\frac{W}{r^{2}}$ (5.5)
where $W$ represents the radial displacement function, $V$ is the perturbation of the gravitational potential, $\omega$ is the angular frequency of oscillations, and $l$ is the spherical harmonic index.
@@
    $\omega^{2}R^{2}e^{\lambda}-2\Phi V+\frac{d\Phi}{dr}W=0$ (5.7)
@@ 6 Relationship between energy density and pressure
6.1 Effective Lagrangian for neutron–dark matter conversion
To model the possibility of neutron decay into dark matter inside neutron stars, we extend the nucleonic sector (described here using the QMC model) by including a dark fermion $\chi$ carrying baryon number. In order to reproduce the repulsive self-interaction assumed in the EoS analysis, we introduce a dark vector mediator $V_{\mu}$, analogous to the $\omega$ meson in relativistic mean-field models. The effective Lagrangian reads
$\mathcal{L}\;=\;\mathcal{L}_{\rm QMC}[n,p,\sigma,\omega_{\mu},\rho_{\mu}]+\bar{\chi}\,(i\not{\partial}-m_{\chi})\,\chi-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+\frac{1}{2}m_{V}^{2}V_{\mu}V^{\mu}-g_{\chi}\,\bar{\chi}\gamma^{\mu}\chi V_{\mu}+\varepsilon\,(\bar{n}\chi+\bar{\chi}n),$ (6.1)
where $V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}$. The last term represents an effective in-medium mixing between neutrons and dark fermions, which allows for neutron conversion into dark matter when $\mu_{n}>\mu_{\chi}$. Baryon number is conserved if $\chi$ carries baryon number =1=1.
@@ 6.1 Effective Lagrangian for neutron–dark matter conversion
$\mathcal{L}_{\rm eff}\supset\frac{g_{\chi}^{2}}{2m_{V}^{2}}\,(\bar{\chi}\gamma_{\mu}\chi)(\bar{\chi}\gamma^{\mu}\chi)\equiv\frac{G}{2}\,(\bar{\chi}\gamma_{\mu}\chi)^{2},$ (6.2)
@@
The results indicate that the nucleon-only EoS is the stiffest among the considered models. As the DM–DM self-repulsion increases, the corresponding EoS for neutron decay into dark matter becomes progressively stiffer and approaches the nucleon-only case which is in agreement with the findings of [50, 41, 51]. This behavior arises because stronger DM–DM self-interaction raises the energy cost for neutron decay into dark matter, thereby reducing the overall dark matter population within the system [50]. Consequently, larger self-interaction strengths, $G$, result in fewer dark matter particles and stiffer equation of state [50].
@@
Figure 2: Mass–radius relations for neutron stars with different equations of state (EoSs). Shown are nucleon-only (QMC), strange matter (MIT bag model), and dark-matter–admixed stars with varying self-interaction strengths $G$. The horizontal band indicates the observed $2,M_{\odot}$ pulsar constraint. Increasing $G$ stiffens the dark-matter EoS, raising the maximum mass toward the nucleon-only case. The radius of light neutron stars is smaller because the EoS is not optimized in the lower energy density region (outer crust region) [109].
@@ 7.1 Damping time
$\dot{E}{\rm{}_{GW}}=-\frac{32G}{5c^{5}}\omega^{6}\left|\int\rho(r)r^{2}Y{2m}(\theta,\phi)drd\Omega\right|^{2}.$ (7.3)
From these relations, the universal scaling $\tau\sim R^{4}/M^{3}$ emerges [125], which we adopt and extend in this study.
@@
Figure 3: The normalised damping time of the f-mode as a function of stellar mass. Nucleon-only stars exhibit the shortest damping times, while strange-matter stars show longer-lived oscillations. Dark-matter–admixed stars interpolate between these limits depending on $G$.
@@ 7.2 Mass-scaled angular frequency
Figures (4) and (5) illustrate the relation between the mass-scaled angular frequency, where $\omega_{r}=2\pi f$, and two structural parameters of neutron stars: compactness ($M/R$) and $\psi=\sqrt{M^{3}/I}$ [117]. These diagnostics are particularly valuable since they encode information about the distribution of mass and the role of internal composition in determining stellar oscillation properties. The parameter $\sqrt{M^{3}/I}$, in particular, is closely tied to the efficiency of gravitational-wave (GW) emission, as it relates the global stellar mass to the moment of inertia, thereby providing insight into how matter is distributed inside the star.
@@
Figure (5) further reinforces the presence of a quasi-universal relation across different microscopic models when the angular frequency is expressed in terms of $\psi=\sqrt{M^{3}/I}$. Remarkably, even when exotic compositions are considered—including strange matter and neutron-decay into DM with strong self-interaction ($G>26$ fm2)—the data points collapse onto a nearly single curve. This indicates that $\psi$ effectively captures the essential structural scaling of neutron stars, smoothing out differences in composition. Deviations are observed only in the dark matter without self-repulsion case, but the overall correlation persists. This reinforces the robustness of universal relations in gravitational-wave asteroseismology.
@@ 8 F-mode frequency
The frequency of an f-mode is highly sensitive to the EoS of the ultra-dense matter within the neutron star [116, 117, 118]. A stiffer EoS generally predicts higher f-mode frequencies. Conversely, a softer EoS predicts lower f-mode frequencies. The acceptable values of f-mode frequency for compact stars, such as neutron stars, typically range between 1 and 1–3 kHz. These frequencies are significant because they fall within the detection range of advanced gravitational wave detectors like the Einstein Telescope and Cosmic Explorer [24]. Figure (6) illustrates that the f-mode frequency of the selected models adheres to the universal relation, demonstrating consistency across different astrophysical scenarios. Notably, both frequency ranges fall within the detection capabilities of the Einstein Telescope and Cosmic Explorer, highlighting their potential for observational verification. Furthermore, for neutron decay into dark matter, variations in dark matter self-interaction strength do not disrupt the universal trend. Regardless of the DM-DM self-interaction parameter, the f-mode frequency remains well within the expected range, reinforcing the robustness of the underlying physical principles governing these oscillations. This consistency suggests that such models could be effectively probed through next-generation gravitational wave observatories.
Our results are consistent with earlier works exploring neutron-to-dark matter decay in neutron stars. Motta et al. [42] and Grinstein et al. [17] found that non-interacting dark fermions destabilize neutron stars, reducing their maximum mass below the $2M_{\odot}$ threshold, which we confirm in our analysis. Husain et al. [50] similarly emphasized the necessity of DM self-repulsion, consistent with our finding that $G>26$ fm2 restores compatibility with observational constraints.
For f-mode oscillations, our study extends the universal relations established by Andersson and Kokkotas [125] and refined by Lioutas and Stergioulas [126], showing that these relations persist even in the presence of exotic matter. However, we demonstrate that strange matter and non-self-interacting DM produce systematic deviations, suggesting that precise future detections could be used to discriminate between EoSs.

### v3

# F-mode Oscillations of Neutron Stars with Dark Matter from Neutron Decay: Implications for Gravitational-Wave Detectability

@@ Abstract
Importantly, it has been shown that for a GW energy release of $E\sim 10^{52}$ erg and a source distance of 25 Mpc, the characteristic strain and signal-to-noise ratio exceed the ET-D sensitivity threshold below $\sim$2.1 kHz for all models except the non-interacting DM case, demonstrating that neutron-to-dark matter decay scenarios, including the role of DM self-interactions, can be tested through next-generation gravitational-wave asteroseismology, offering a new probe of DM physics and the neutron lifetime anomaly.
keywords:
Neutron stars , Dark Matter , Neutron decay , f-mode oscillations.
journal: Physics of Dark Universe
[…]
organization=School of Science, Technology and Engineering, University of Sunshine Coast, ,addressline=182, Victoria Square, city=Adelaide, postcode=5000, state=SA, country=Australia
@@ 1 Introduction
The fundamental mode (f-mode) is a type of non-radial oscillation in compact stars, primarily governed by the star’s bulk properties such as mass, radius, and internal composition. These oscillations generate gravitational waves, and their frequency and damping time are sensitive to the equation of state (EoS) of the stellar interior. As such, the f-mode serves as a powerful diagnostic tool for probing the internal structure and composition of neutron stars and, more generally, for constraining the nature of dense matter under extreme conditions.
@@
Furthermore, the damping time and frequency of the f-mode are also linked to the star’s properties, such as mass, compactness, and moment of inertia, quantities that could be inferred from spin-orbit coupling effects in binary pulsar systems [27]. These interdependencies allow the f-mode to act as a bridge between observable gravitational wave signatures and the otherwise inaccessible interior physics of compact stars. In this work, previous studies of neutron decay into dark matter have been extended by incorporating both nucleonic and strange matter EoS within the quark–meson coupling (QMC) model, and by explicitly evaluating the detectability of f-mode oscillations with next-generation gravitational-wave detectors. While earlier works have addressed stability and mass-radius constraints for such stars, a systematic investigation of characteristic strain and signal-to-noise ratios (SNR) in the context of ET-D sensitivity has not yet been carried out. These results therefore provide the first quantitative assessment of whether gravitational-wave observations of f-modes can constrain dark matter self-interactions and the presence of exotic matter in neutron stars.
This paper presents, for the first time, a systematic test of the neutron–to–dark matter decay model using f-mode oscillations and signal-to-noise ratio estimates. The study is organized as follows. Section 2 introduces the theoretical framework for neutron decay into dark matter and discusses astrophysical constraints. Section 3 provides an overview of neutron star formation, structure, and observed phenomena, while Section 4 addresses the equations of state governing ultradense matter. Section 5 outlines the stellar structure equations, and Section 6 develops advanced considerations of the EoS. Section 7 examines universal relations of f-mode oscillations, with subsections devoted to damping times (7.1) and mass-scaled angular frequencies (7.2). Section 8 compares f-mode frequencies predicted by nucleonic, strange matter, and dark matter models. Section 9 evaluates detectability by confronting the predicted f-mode signals with the Einstein Telescope sensitivity curve and presenting signal-to-noise ratios. Section 10 concludes with a summary of the main results and their implications for neutron star physics, gravitational-wave astronomy, and the search for dark matter signatures.
@@ 4 Equation of State
$\epsilon_{H}=\frac{1}{2}m^{2}_{\sigma}\sigma^{2}+\frac{1}{2}m^{2}_{\omega}\omega^{2}+\frac{1}{2}m^{2}_{\rho}\rho^{2}+\frac{1}{\pi^{2}}\int^{k_{F}^{n}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk+$ (2)
$\frac{1}{\pi^{2}}\int^{k_{F}^{p}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{e}}_{0}k^{2}\sqrt{k^{2}+m_{e}^{2}}dk+$
$\frac{1}{\pi^{2}}\int^{k_{F}^{\mu}}_{0}k^{2}\sqrt{k^{2}+m_{\mu}^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{\chi}}_{0}k^{2}\sqrt{k^{2}+m_{\chi}^{2}}dk$
where $m_{i}$ is the mass of the corresponding particle (subscripted), and the details of the last term are given in section  6.1. The nucleon effective mass, $M_{N}^{*}(\sigma)$, given by [106, 107]
$M_{N}^{*}(\sigma)=M_{N}-g_{\sigma}\sigma+\frac{d\tilde{(}g_{\sigma}\sigma)^{2}}{2}\,.$ (3)
Here $g_{\sigma}$ is the $\sigma$-nucleon coupling constant in free space calculated in the model and $d$ is the scalar polarizability parameter of the nucleon in the QMC model, which accounts for quadratic corrections to the scalar field coupling [106]. The pressure ($P$) is calculated using
$P=\int_{i}\mu_{i}n_{i}-\varepsilon,$ (4)
where $i=n,p,e,\mu,\chi$. The nucleons-only EoS is taken from [106] and strange matter EoS is drive based on MIT bag model, and the EoS used is given by
$P=\frac{1}{3}(\varepsilon-4B),$ (5)
[…]
@@ 5 Structural equations
Neutron stars are dense astrophysical objects shaped by intense nuclear and gravitational forces. The potential presence of dark matter within these stars may influence their structural characteristics, resulting in changes to their equation of state (EoS). This study examines methods for integrating dark neutron components into the conventional neutron star EoS and solving the Tolman-Oppenheimer-Volkoff (TOV) [112, 113] equations to assess their impact.
@@
$\frac{dP}{dr}=-\frac{(\epsilon+P)}{r}\frac{d\Phi}{dr}$ (6)
$\frac{d\Phi}{dr}=\frac{M+4\pi r^{3}P}{r(r-2M)}$ (7)
$\frac{dM}{dr}=4\pi r^{2}\epsilon$ (8)
where $P$ is the pressure, $\epsilon$ is the energy ensity, $\Phi$ is the gravitational potential, $M$ is the mass, and $r$ is the radial coordinate.
@@
The Lindblom–Detweiler formalism for non-radial oscillations [114] has been adopted in this study. Here $\nu(r)$ and $\lambda(r)$ are the static metric functions in the line element $ds^{2}=-e^{\nu}dt^{2}+e^{\lambda}dr^{2}+r^{2}d\Omega^{2}$, and $(X,W,V)$ are perturbation variables defined as in [115]. This ensures dimensional consistency and reproducibility.
The f-mode oscillations are governed by two coupled differential equations [116, 117, 118]
$\frac{dX}{dr}=-\frac{l+1}{r}X+\frac{e^{\lambda/2}}{r}\left[\omega^{2}re^{-\nu/2}V+\frac{d\nu}{dr}W\right],$ (9)
$\frac{dW}{dr}=-\frac{d\nu}{dr}W-\frac{l(l+1)}{r}V+e^{\lambda/2}X,$ (10)
where $W(r)$ represents the radial displacement function, $V(r)$ is the metric perturbation function, $\omega$ is the angular oscillation frequency, and $l$ is the spherical harmonic index of the perturbation.
@@
    $\omega^{2}R^{2}e^{\lambda}-2\Phi V+\frac{d\Phi}{dr}W=0$ (12)
@@ 6 Relationship between energy density and pressure
6.1 Effective Lagrangian for neutron to dark matter decay
To model the possibility of neutron decay into dark matter inside neutron stars, The nucleonic sector (described here using the QMC model) has been extended by including a dark fermion $\chi$ carrying baryon number. In order to reproduce the repulsive self-interaction assumed in the EoS analysis, a dark vector mediator $V_{\mu}$ has been introduced, analogous to the $\omega$ meson in relativistic mean-field models. The effective Lagrangian reads
$\mathcal{L}\;=\;\mathcal{L}_{\rm QMC}[n,p,\sigma,\omega_{\mu},\rho_{\mu}]+\bar{\chi}\,(i\not{\partial}-m_{\chi})\,\chi-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+\frac{1}{2}m_{V}^{2}V_{\mu}V^{\mu}-g_{\chi}\,\bar{\chi}\gamma^{\mu}\chi V_{\mu}+\varepsilon\,(\bar{n}\chi+\bar{\chi}n),$ (13)
where $V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}$. The last term represents an effective in-medium mixing between neutrons and dark fermions, which allows for neutron conversion into dark matter when $\mu_{n}>\mu_{\chi}$. Although the effective in-medium operator $\varepsilon\bar{n}\chi+\text{h.c.}$ mixes neutrons with dark fermions, baryon number is conserved if the dark fermion $\chi$ carries baryon number $B(\chi)=1$. In this case, the conversion $n\leftrightarrow\chi$ redistributes baryon number between visible and dark sectors but does not violate overall conservation. This treatment follows the effective approach proposed in [16, 18, 17].
@@ 6.1 Effective Lagrangian for neutron to dark matter decay
$\mathcal{L}_{\rm eff}\supset\frac{g_{\chi}^{2}}{2m_{V}^{2}}\,(\bar{\chi}\gamma_{\mu}\chi)(\bar{\chi}\gamma^{\mu}\chi)\equiv\frac{G}{2}\,(\bar{\chi}\gamma_{\mu}\chi)^{2},$ (14)
@@
The results indicate that the nucleon-only EoS is the stiffest among the considered models. As the DM–DM self-repulsion increases, the corresponding EoS for neutron decay into dark matter becomes progressively stiffer and approaches the nucleon-only case, which is in agreement with the findings of [48, 39, 49]. This behavior arises because stronger DM–DM self-interaction raises the energy cost for neutron decay into dark matter, thereby reducing the overall dark matter population within the system [48]. Consequently, larger self-interaction strengths, $G$, result in fewer dark matter particles and stiffer equation of state [48].
@@
Figure 2: Mass–radius relations for neutron stars with different equations of state (EoSs). Shown are nucleon-only (QMC), strange matter (MIT bag model), and dark-matter–admixed stars with varying self-interaction strengths $G$. The radius of light neutron stars is smaller because the EoS is not optimized in the lower energy density region (outer crust region) [107].
@@ 7.1 Damping time
$\dot{E}{\rm{}_{GW}}=-\frac{32G}{5c^{5}}\omega^{6}\left|\int\rho(r)r^{2}Y{2m}(\theta,\phi)drd\Omega\right|^{2}.$ (17)
From these relations, the universal scaling $\tau\sim R^{4}/M^{3}$ emerges [125], which has been adopted and extended in this study.
@@
Figure 3: The normalised damping time of the f-mode as a function of stellar mass. Nucleon-only stars exhibit the shortest damping times, while strange-matter stars show longer-lived oscillations.
@@ 7.2 Mass-scaled angular frequency
Figures (4) and (5) illustrate the relation between the mass-scaled angular frequency, where $\omega_{r}=2\pi f$, and two structural parameters of neutron stars, namely, compactness ($M/R$) and $\psi=\sqrt{M^{3}/I}$ [117]. These diagnostics are particularly valuable since they encode information about the distribution of mass and the role of internal composition in determining stellar oscillation properties. The parameter $\sqrt{M^{3}/I}$, in particular, is closely tied to the efficiency of gravitational-wave (GW) emission, as it relates the global stellar mass to the moment of inertia, thereby providing insight into how matter is distributed inside the star.
@@
Figure (5) further reinforces the presence of a quasi-universal relation across different microscopic models when the angular frequency is expressed in terms of $\psi=\sqrt{M^{3}/I}$. Remarkably, even when exotic compositions are considered, including strange matter and neutron-decay into DM with strong self-interaction ($G>26$ fm2), the data points collapse onto a nearly single curve. This indicates that $\psi$ effectively captures the essential structural scaling of neutron stars, smoothing out differences in composition. Deviations are observed only in the dark matter without self-repulsion case, but the overall correlation persists. This reinforces the robustness of universal relations in gravitational-wave asteroseismology.
@@ 8 F-mode frequency
The frequency of an f-mode is highly sensitive to the EoS of the ultra-dense matter within the neutron star [116, 117, 118]. A stiffer EoS generally predicts higher f-mode frequencies. Conversely, a softer EoS predicts lower f-mode frequencies. The acceptable values of f-mode frequency for compact stars, such as neutron stars, typically range between 1 and 1–3 kHz. These frequencies are significant because they fall within the detection range of advanced gravitational wave detectors like the Einstein Telescope and Cosmic Explorer [24]. Figure (6) illustrates that the f-mode frequency of the selected models adheres to the universal relation, demonstrating consistency across different astrophysical scenarios.
Notably, both frequency ranges fall within the detection capabilities of the Einstein Telescope and Cosmic Explorer, highlighting their potential for observational verification. Furthermore, for neutron decay into dark matter, variations in dark matter self-interaction strength do not disrupt the universal trend. Regardless of the DM-DM self-interaction parameter, the f-mode frequency remains well within the expected range, reinforcing the robustness of the underlying physical principles governing these oscillations. This consistency suggests that such models could be effectively probed through next-generation gravitational wave observatories.
For f-mode oscillations, this study extends the universal relations established by Andersson and Kokkotas [125] and refined by Lioutas and Stergioulas [127], showing that these relations persist even in the presence of exotic matter. However, it has been demonstrated that strange matter and non-self-interacting DM produce systematic deviations, suggesting that precise future detections could be used to discriminate between EoSs.
@@
Figure 7 presents the predicted characteristic strain of f-mode oscillations for the considered EoS models, compared with the ET-D sensitivity curve [129, 130]. The strain values have been calculated assuming a fiducial gravitational-wave energy release of E = 1052 erg, a source distance of D = 25Mpc, and a damping time of $\tau$ = 0.1s, consistent with estimates in [131, 132]. Under these benchmark conditions, the non–self-interacting dark matter EoS (G0) yields strain amplitudes that remain below the ET sensitivity threshold across the relevant frequency range, implying that such signals would not be detectable. By contrast, the self-interacting dark matter and strange matter EoSs predict higher characteristic strains that exceed the ET sensitivity curve for frequencies below 2.1 kHz. This demonstrates that, if sources within 25 Mpc deposit of order 1052 erg into f-mode oscillations, the corresponding signals could in principle be detected by ET. These results emphasize that f-mode detectability is highly sensitive to the underlying neutron-star EoS, and they suggest that next-generation detectors such as ET can provide meaningful constraints on dark matter interactions through neutron-star asteroseismology.
Figure 7: Characteristic strain $h_{c}$ of f-mode oscillations from neutron stars as a function of frequency, compared with the sensitivity curve of the Einstein Telescope (ET-D) [130]. Results are shown for several equations of state (EoS), assuming a total gravitational-wave energy release of E = 1052 erg, damping time = 0.1 sec and a source distance of D = 25 Mpc.
Figure (8) shows the variation of SNR with f-mode frequency for the different EoS models considered. It has been found that the SNR exceeds unity for frequencies below approximately 2.1 kHz. This implies that, within this frequency range, f-mode oscillations could be detectable by the Einstein Telescope, provided the assumed energy budget and source distance. At higher frequencies, the SNR falls below unity, indicating that detection becomes unlikely due to the reduced strain amplitude and the rising noise floor of the detector. Notably, the non–self-interacting DM model (G0) yields systematically lower SNRs, while the interacting dark matter and strange matter EoS predict higher SNR values, suggesting more favorable detectability. These results emphasize that the detection prospects of f-mode signals strongly depend on both the EoS and the frequency band of emission, with ET’s sensitivity window favoring signals in the 2.1 kHz regime.
Figure 8: Signal-to-noise ratio (SNR) of f-mode gravitational-wave signals as a function of frequency, computed for different equations of state (EoS) at a source distance of D = 25 Mpc, damping time = 0.1 sec and energy release E = 1052 erg.
The detectability of f-mode gravitational-wave signals is highly sensitive to both the neutron star equation of state (EoS) and the frequency band in which the mode energy is emitted.
The physics behind these trends can be understood as follows. Softer equations of state, such as those with exotic matter components, generally produce lower f-mode frequencies, shifting the emission closer to the frequency band where ET is most sensitive. Moreover, longer damping times associated with strange matter enhance the overall energy radiated in the mode, further increasing the characteristic strain. Conversely, the absence of DM self-repulsion (G0) leads to reduced pressure support, suppressing the excitation energy available to the mode and lowering the strain amplitude. Thus, the inclusion of dark matter with repulsive self-interactions not only restores astrophysical stability but also improves the prospects for gravitational-wave detection.
From an observational standpoint, these results indicate that ET has the potential to place meaningful constraints on the nature of dark matter interactions and the composition of neutron stars. A confirmed detection of f-mode oscillations in the $\sim$1–2 kHz range would favor scenarios involving exotic matter or self-interacting dark matter, whereas a systematic absence of such signals could be used to rule out or constrain these models. Future work should extend this analysis to population studies, varying source distances, masses, and energy release efficiencies, in order to more comprehensively assess the detectability of f-modes in realistic astrophysical environments.